OpenStudy (anonymous):

What is the simplest form of the expression? (2√−6√)over(2√+6√)

OpenStudy (ybarrap):

Is this the problem? $$ \cfrac{2\sqrt{-6}}{2\sqrt{6}}\\ =\cfrac{\sqrt{-6}}{\sqrt{6}}\\ =\sqrt{\cfrac{-6}{6}}\\ =\sqrt{-1}\\ =i $$

OpenStudy (anonymous):

\[\sqrt{2}-\sqrt{6}\over \sqrt{2}+\sqrt{6}\]

OpenStudy (anonymous):

@ybarrap sorry thats the equation

OpenStudy (ybarrap):

$$ {\sqrt{2}-\sqrt{6}\over \sqrt{2}+\sqrt{6}}\\ ={\sqrt{2}-\sqrt{2\times3}\over \sqrt{2}+\sqrt{2\times3}}\\ ={\sqrt{2}\left( 1-\sqrt{3}\right )\over \sqrt{2}\left (1+\sqrt{3}\right )}\\ ={\left( 1-\sqrt{3}\right )\over \left (1+\sqrt{3}\right )}\\ ={\left( 1-\sqrt{3}\right )\over \left (1+\sqrt{3}\right )}\times\cfrac{1-\sqrt{3}}{1-\sqrt{3}}\\ =\cfrac{1-2\sqrt 3+3}{1-3}\\ =\cfrac{4-2\sqrt 3}{-2}\\ =-2+\sqrt 3\\ $$

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

@ybarrap : check last step, sign mistake. Medal n Fan

OpenStudy (anonymous):

ucan go tohttp:://wolframalpha.com u will get steps and graphs along with answer

OpenStudy (anonymous):

@anubhav.m is your answer correct?

OpenStudy (anonymous):

sorry for the typos

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

It should be.....y? Is it wrong?

OpenStudy (anonymous):

yep, it correct....

OpenStudy (anonymous):

@ybarrap got the same answer......sry abt that, there's no mistake.....